OpenStudy (anonymous):

find the lines that are tangent and normal to the curve at the given point x^2y^2=81 at the point (1,9)

OpenStudy (anonymous):

to get the tangent line you have to find its slope and a point now you have the point (1,9) how to get the slope...any ideas ?

OpenStudy (anonymous):

solve for x im guessing

OpenStudy (anonymous):

look The first derivative means the slope of the tangent

OpenStudy (anonymous):

would the derivative be 2x2y

OpenStudy (anonymous):

\[x ^{2}*y ^{2}=81\] right form ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ok how the result is 2x*2y ?

OpenStudy (anonymous):

idk what to do

OpenStudy (anonymous):

\[\frac{ d }{ dx }(x ^{2}y ^{2})=\frac{ d }{ dx }(81)\]

OpenStudy (anonymous):

\[2x*y ^{2}+2yx ^{2}*\frac{ dy }{ dx }=0\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ -y }{ x }\]

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

do you plug the point values in?

OpenStudy (anonymous):

yes go ahead

OpenStudy (anonymous):

dy/dx=?

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

good so the slope is -9 and the point is (1,9) now can you create the tangent line ?

OpenStudy (anonymous):

y=-9x+18

OpenStudy (anonymous):

how would you find the line normal to the curve

OpenStudy (anonymous):

the line normal to the curve at a point is also normal to the tangent line at the same point so the slope of the normal line=-1/m where m=slope of tangent line so the slope of the normal line =??

OpenStudy (anonymous):

so just the inverse which is -1/9

OpenStudy (anonymous):

no 1/9 not -1/9

OpenStudy (anonymous):

what would the equation be i got y=1/9x+82/9

OpenStudy (anonymous):

80/9 not 82/9

OpenStudy (anonymous):

cool thanks so much

OpenStudy (anonymous):

wait check this :) the blue is your function the red line is the tangent and the green is the normal at the same point